### v2

# New Encoders for German Trained from Scratch: Comparing ModernGBERT with Converted LLM2Vec Models

Abstract
Encoders remain essential for efficient German NLP and NLU scenarios despite the rise of decoder-only LLMs.
This work studies two routes to high-quality German encoders under identical data and training constraints:
a) training from scratch and
b) converting decoders via LLM2Vec.
We introduce two resources:
ModernGBERT (134M, 1B), fully transparent German encoders in the ModernBERT style, and
LLäMmlein2Vec (120M, 1B, 7B), decoder-to-encoder conversions trained with masked next-token prediction,
both undergoing a context extension to tokens.
Across SuperGLEBer, ModernGBERT 1B sets a new state of the art (avg 0.808), surpassing GBERTLarge (+4%) and the seven-times larger converted 7B model (0.787).
On German MTEB after supervised fine-tuning, ModernGBERT 1B (0.551) approaches the converted 7B model (0.557).
We release all models, checkpoints, datasets, and full training records, and introduce an encoder-adapted QA-NIAH evaluation.
All in all, our results provide actionable guidance:
when parameter efficiency and latency matter, from-scratch encoders dominate. When a pre-trained decoder exists and compute is a limited, conversion offers an effective alternative.111ModernGBERT and LLäMmlein2Vec, including all code, data and intermediary checkpoints will be published upon acceptance under a research-only RAIL license.
Keywords: Language Resources, Encoder-only Models, German, Model Conversion, Long Context, Benchmarking
New Encoders for German Trained from Scratch:
Comparing ModernGBERT with Converted LLM2Vec Models
| Julia Wunderle1,∗, Anton Ehrmanntraut2,∗, Jan Pfister1 |
| Fotis Jannidis2,†, Andreas Hotho1,† |
| 1Data Science 2Computer Philology and History of Contemporary German Literature |
| CAIDAS – Center for Artificial Intelligence and Data Science |
| JMU – Julius-Maximilians-Universität Würzburg |
| {lastname}@informatik.uni-wuerzburg.de, {firstname.lastname}@uni-wuerzburg.de |
| Equal contribution |
Abstract content
1. Introduction
Encoders remain central to efficient (German) NLP and NLU where latency, memory, and cost dominate Pfister and Hotho (2024), still accounting for most of the downloads from e.g. Hugging Face Bourdois (2025). Especially for tasks like sentence similarity (i.e. important for RAG), mask-filling and text/token classification, encoder models are still the architecture of choice Bourdois (2025). Yet, the German NLP landscape lacks high-quality, modern, transparent encoders with reproducible training provenance and long-context support. This paper contributes both resources and a controlled study addressing two practical routes to encoders under identical data and training constraints: training from scratch versus converting decoders.
We introduce to complementary model families: ModernGBERT (134M, 1B), ModernBERT-style German encoders trained from scratch, and LLäMmlein2Vec (120M, 1B, 7B), encoders derived from German decoder-only models via LLM2Vec. ModernBERT (Warner et al., 2024) introduced several architectural improvements for English encoders, including improved relative positional embeddings and efficient attention patterns that allow long context processing. We adapt this design to German, providing strong encoder baselines built entirely from scratch.
In parallel, to assess the practical utility and trade-offs of training encoder models from scratch, we converted the decoder-only model family LLäMmlein into encoders using LLM2Vec (BehnamGhader et al., 2024). Specifically, to align closely with the encoder training objective, we limit the procedure to the first two LLM2Vec steps: 1. enabling a bidirectional attention mask, and 2. masked next-token prediction (MNTP). Since all models share the same training datasets, this setup provides a unique foundation for systematically analyzing the relationship between different architectures and training strategies. We extensively evaluate and compare these models during and post training via: natural language understanding (SuperGLEBer, Pfister and Hotho, 2024), embedding performance (MTEB; Enevoldsen et al., 2025; Muennighoff et al., 2023; Wehrli et al., 2023), long-context understanding (new Question Answering Needle-in-a-Haystack (QA-NIAH)) and an efficiency suite reflecting variable-length inference. Our key contributions are:
We introduce a ModernGBERT family, which achieves new state-of-the-art performance on SuperGLEBer and the German MTEB.
To enable comparisons between training strategies, we also introduce a decoder-turned-encoder LLäMmlein2Vec family, based on the same training dataset.
We find that dedicated encoders trained from scratch consistently outperform converted decoders of similar size.
Our newly introduced QA-NIAH benchmark confirms strong long-context understanding of our models, validating the effectiveness for extended input sequences.
To enable further research we release all (intermediary) resources including all model checkpoints, data point tracking, code and data.
2. Datasets
2.1. Pre-Training Dataset
We pre-trained ModernGBERT on the same data as LLäMmlein decoder models (Pfister et al., 2024), using the open-source RedPajamaV2 dataset Weber et al. (2024) . This dataset comprises German CommonCrawl snapshots from 2014 to 2023. As we intend to keep datasets constant between ModernGBERT and LLäMmlein, we follow LLäMmlein’s data pipeline and select the higher quality document-level deduplicated “head” and “middle” partitions, excluding the lower quality “tail” partition. For our 134M model, we only selected the head partition. Our processing pipeline mirrors Pfister et al. (2024): we first perform paragraph-level deduplication using a Bloom filter to remove redundant content such as GDPR notices and web boilerplate, enhancing data diversity. Afterward, we apply a token-to-word ratio filter to further improve text quality. The final dataset is about , corresponding to approximately 1.27T tokens using a GBERTLarge tokenizer.
2.2. Context Extension Datasets
ModernBERT enhances its context capacity from to by fine-tuning in two phases: on an 250B-token subset of -token sequences from the original pre-training data (ext1), followed by 50B-token high-quality dataset with mixed sequence lengths (ext2) Gao et al. (2025).
Following this setup, we construct two German datasets. For ext1, we analogously subsample long sequences from our pre-training dataset, resulting in “LONG-Head” from the head partition for our 134M model, and “LONG-Head/Middle” from the head and middle partition for our 1B model.
For the high-quality dataset used for ext2, we selected the German portion of Fineweb2 as basis (Penedo et al., 2024) . To match the original distribution, we take a randomized sample of Fineweb2, and add a separate Fineweb2 sample, selecting long documents with tokens, splitting them into sequences of about tokens (“Fineweb2-long”). Additional long documents are drawn from the 2023 German Wikipedia and the 2022 OpenLegalData dump, . The resulting dataset contains 14.4B tokens, summarized in Table˜1.
| Dataset | # Tokens | # Sequences | Median Length | |
ext1 |
LONG-Head | B | ||
| LONG-Head/Middle | B | |||
ext2 |
High Quality | B | ||
M | |||
M | |||
M | |||
M |
3. Methodology 111All training settings and parameters will be detailed in the camera-ready version.
3.1. ModernGBERT
We adapt the ModernBERT architecture and training strategy for German. While ModernGBERT 134M matches the base ModernBERT model size (22 layers, 768 hidden units, but 16M fewer parameters due to a smaller vocabulary size), we create ModernGBERT 1B with 28 layers and a hidden size of .
Pre-/Post-Training Strategy
Both models are pre-trained using masked language modeling (MLM) with no next-sentence-prediction, a 30% masking rate, and sequences up to tokens ( RoPE theta). ModernGBERT 134M is trained on the head partition (0.47T tokens), as downstream evaluation indicated early saturation, while ModernGBERT 1B is trained on both the head and middle partitions of our pre-training corpus (Section˜2.1), totaling 1.27T tokens.
After standard MLM pre-training, we proceed with the two context extension phases: During ext1, the RoPE theta is raised to , and models are trained on the LONG-Head (134M) or LONG-Head/Middle (1B) datasets (see Section˜2.2). In the ext2 phase, both models are trained on the High Quality dataset.
Tokenization
For tokenization, we use the original BERT-style tokenizer from GBERTLarge, resulting in a -word embedding layer. While LLäMmlein (Pfister et al., 2024) provides a dedicated German BPE tokenizer, our preliminary ablations using this tokenizer consistently showed degraded downstream performance. Consequently, we retained the GBERTLarge tokenizer.
Training Progress Tracking
Throughout the training we save, evaluate and release all checkpoints to support further research. Inspired by Pythia (Biderman et al., 2023) we provide full training transparency by logging and releasing the order of data points seen during training; thus, all checkpoints can be linked with the exact data points seen up to that checkpoint.
3.2. LLM2Vec: Turning Decoders to Encoders
BehnamGhader et al. (2024) proposes a method to convert decoder-only LLMs into text encoders through the following steps: First, the causal attention mask is replaced with a full attention mask, enabling bidirectional attention across tokens. Second, the model is trained using a masked next token prediction (MNTP) objective. Third, regular LLM2Vec includes a unsupervised contrastive learning (SimCSE) step, improving embedding quality by maximizing agreement between differently dropped-out versions of the same input. However, we intend to remain as close as possible to the training objectives of ModernGBERT to allow fair and direct comparisons. Therefore, we limit the procedure to the first two steps of LLM2Vec, specifically employing the MNTP objective, which is most closely aligned with MLM.
Pre-/Post-Training Strategy
We train all three LLäMmlein models (120M, 1B, 7B) using the same two context extension datasets as employed by ModernGBERT’s context extensions (ext1 and ext2) (Section˜2.2). Specifically, the 120M model mirrors ModernGBERT 134M, being trained on the LONG-Head (ext1) and High Quality (ext2) datasets, while the 1B and 7B models follow ModernGBERT 1B, using LONG-Head/Middle (ext1) and the High Quality dataset (ext2). Notably, MNTP training is applied separately to each dataset, resulting in two adapter modules for each model. Additionally, all adapters were trained on the full corresponding datasets333except for the LLäMmlein2Vec 7B trained on the LONG-Head/Middle model, which we trained on 64 nodes with 4 H200 each for 14 hours, before stopping the training due to compute constraints., although models were able to achieve comparable results during earlier training stages. This indicates, that compute time could have been drastically reduced.
Further aligning with ModernGBERT we combined MNTP with context extension — a setup that, to our knowledge, has not yet been widely explored. Therefore, we also extended the model sequence length to tokens and increased RoPE theta to . We evaluate both individual adapters (ext1 & ext2) as well as a merged model (ext1+2) that combines both
4. Evaluation Setup
4.1. SuperGLEBer
We assess our final models using the German SuperGLEBer benchmark Pfister and Hotho (2024), which includes 29 tasks across text classification, sequence tagging, question answering, and sentence similarity. These tasks cover diverse domains such as news, legal texts, and consumer reviews. For each task, models are fine-tuned with QLoRA Dettmers et al. (2023) by default, or LoRA as fallback. In addition to evaluating final checkpoints, we follow LLäMmlein Pfister et al. (2024) and evaluate intermediate checkpoints of ModernGBERT as well as LLäMmlein2Vec on the same representative SuperGLEBer subset consisting of: the classification tasks NLI Conneau et al. (2018), FactClaiming Comments Risch et al. (2021), DB Aspect Wojatzki et al. (2017), and WebCAGe Henrich et al. (2012), the sequence tagging task EuroParl Faruqui and Padó (2010), and the sentence similarity task PAWSX Liang et al. (2020).
4.2. Massive Text Embedding Benchmark
We further evaluate the models on the German subset of the Massive Text Embedding Benchmark MTEB(deu,v1) (Enevoldsen et al., 2025), comprising 19 tasks. In addition to text pair classification and semantic textual similarity, which are already covered by the SuperGLEBer benchmark, MTEB includes clustering Wehrli et al. (2023), as well as reranking and retrieval tasks. These latter tasks provide a more comprehensive assessment of general-purpose sentence embeddings, focusing on the models’ ability to produce robust semantic representations.
To adapt the base models for embedding tasks, we fine-tune them using the Sentence-Transformer framework (Reimers and Gurevych, 2019) in a supervised setup. Fine-tuning employs samples from the German portion of the machine-translated multilingual mMARCO passage ranking dataset (Bonifacio et al., 2022), maximizing similarity between query and positive passages, while minimizing similarity to negative passages. Sentence embeddings are obtained by mean pooling over the final token representations. We use InfoNCE loss with a batch size of 128 and a learning rate of . We apply QLoRA for efficient training (falling back to LoRA for the GBERT family, where quantization is not supported).
4.3. Long-Context Understanding
Evaluating long-context capabilities in German is hindered by the scarcity of native high-quality datasets, with translations from English often introducing artifacts. To address this, we construct a new Question-Answering Needle-In-a-Haystack (QA-NIAH) evaluation (Ivgi et al., 2023; Hsieh et al., 2024) based on the human-annotated GermanQuAD dataset (Möller et al., 2021). To our knowledge, this is the first QA-NIAH evaluation specifically adapted for encoder models. The goal in QA-NIAH is to extract an answer span from a long document. We adapt GermanQuAD to a QA-NIAH setup as follows: for each question-paragraph (“needle”) pair, we sample up to 3 distractor paragraphs and shuffle them with the needle, forming a “haystack” document of up to tokens. The answer always appears only in the needle paragraph. For evaluation, up to 20 distractors are included to test generalization to longer contexts, yielding documents up to tokens. This results in training and test question–haystack pairs. Models were again fine-tuned using QLoRA falling back to LoRA, where quantization is not supported. We will release the QA-NIAH generation code and splits to facilitate reuse and extension.
5. Evaluation Results
The tables in this paper present only a summary of our results. Extensive benchmarking results, analyses, and all hyperparameters of all models will be provided in the camera-ready version.
5.1. Intermediate Model Evaluation
ModernGBERT Training
To track the pre-training progress, we evaluated intermediate checkpoints on the SuperGLEBer benchmark (see Section˜4.1). Figure˜2 exemplary shows the results for the ModernGBERT 1B model in the top row. To quantify trends, we first assessed the full SuperGLEBer suite on five 1B and respective three 134M equally spaced ModernGBERT checkpoints and performed Wilcoxon signed-rank tests. The 134M model plateaued after 72B tokens (15% of data), with no further significant gains. In contrast, the 1B model achieved significant improvements after the same amount of data (p 0.0001) and on the middle partition (p 0.00052), before plateauing after 864B tokens (67% of data), with only minor gains thereafter (0.777 0.791)
To gain more insights, we tracked six representative SuperGLEBer subtasks at each checkpoint. For the 134M model, only PAWSX showed a positive Spearman rank correlation with training duration (r = 0.655; p 0.003), whereas for 1B, all tasks except EuroParl improved significantly (r 0.57; p < 0.00014), and complex tasks like PAWSX continued to show modest gains even after the aggregate score stabilized (gray in Figure˜2).
These saturation patterns, including per-task trends and overall performance plateaus, are consistent with findings by Pfister et al. (2024) for decoder models, and by Antoun et al. (2024) for the French ModernBERT variant ModernCamemBERT (136M parameters). Our results confirm that while small ModernBERT models saturate quickly, larger models continue to benefit from additional data. Extrapolating from this observed scaling behaviour between ModernGBERT 134M and 1B, we hypothesize that a larger 7B encoder could leverage extensive monolingual corpora to achieve performance beyond ModernGBERT 1B.
LLäMmlein2Vec Conversion
To remain as consistent as possible with the ModernGBERT training procedure, we trained the LLäMmlein2Vec conversions on the same post-training datasets. To assess whether training on the full dataset was necessary, we evaluated multiple checkpoints throughout the entire training process on the six SuperGLEBer tasks. In the bottom half of Figure˜2 we illustrate the average SuperGLEBer performance alongside the of the standard decoder models (blue), as well as the performance of decoders with a bidirectional attention mask (purple). Interestingly, for encoder-typical tasks such as sentence similarity (PAWSX), the performance of both the 120M, 1B and 7B LLäMmlein models increased even without explicit bidirectional continued pre-training, showing gains from bidirectional fine-tuning alone (120M: 0.477 0.516; 1B: 0.548 0.580; 7B: 0.524 0.660).
Regarding the increases in average performance, the scores remained largely consistent over time, without showing significant improvements. This suggests that training could have been stopped earlier, thus reducing the overall training time, emphasizing that converting a decoder is substantially faster than training a model from scratch.
5.2. Final Model Evaluation
| Model | Size | Avg. | Class. | NER | PAWSX | QA |
| GBERTBase | 111M | 0.718 | 0.723 | 0.786 | 0.561 | 0.803 |
| GBERTLarge | 337M | 0.768 | 0.785 | 0.799 | 0.654 | 0.832 |
| GeBERTaBase | 139M | 0.716 | 0.715 | 0.778 | 0.559 | 0.813 |
| GeBERTaLarge | 406M | 0.749 | 0.743 | 0.791 | 0.619 | 0.844 |
| GeBERTaXLarge | 887M | 0.767 | 0.770 | 0.807 | 0.643 | 0.848 |
| XLM-RoBERTaBase | 279M | 0.689 | 0.693 | 0.754 | 0.505 | 0.802 |
| XLM-RoBERTaLarge | 561M | 0.730 | 0.714 | 0.787 | 0.583 | 0.837 |
| XLM-RoBERTaXLarge | 3.48B | 0.758 | 0.750 | 0.802 | 0.656 | 0.822 |
| mmBERT-small | 142M | 0.748 | 0.759 | 0.800 | 0.606 | 0.828 |
| mmBERT-base | 309M | 0.758 | 0.757 | 0.804 | 0.622 | 0.849 |
| EuroBERT-210m | 212M | 0.706 | 0.659 | 0.698 | 0.607 | 0.859 |
| EuroBERT-610m | 609M | 0.721 | 0.633 | 0.750 | 0.633 | 0.869 |
| EuroBERT-2b | 2.11B | 0.731 | 0.633 | 0.764 | 0.655 | 0.873 |
| LLäMmlein | 120M | 0.676 | 0.702 | 0.712 | 0.477 | 0.812 |
| LLäMmlein2Vec (ext1) | 120M | 0.684 | 0.703 | 0.741 | 0.472 | 0.819 |
| LLäMmlein | 1B | 0.733 | 0.781 | 0.773 | 0.548 | 0.828 |
| LLäMmlein2Vec (ext1) | 1B | 0.762 | 0.776 | 0.812 | 0.615 | 0.843 |
| LLäMmlein | 7B | 0.747 | 0.810 | 0.805 | 0.524 | 0.851 |
| LLäMmlein2Vec (ext1) | 7B | 0.787 | 0.799 | 0.838 | 0.670 | 0.842 |
| ModernGBERT | 134M | 0.730 | 0.716 | 0.782 | 0.589 | 0.833 |
| ModernGBERT (ext1+2) | 134M | 0.749 | 0.735 | 0.805 | 0.612 | 0.836 |
| ModernGBERT | 1B | 0.800 | 0.806 | 0.839 | 0.681 | 0.874 |
| ModernGBERT (ext1+2) | 1B | 0.808 | 0.812 | 0.845 | 0.699 | 0.876 |
Natural Language Understanding
We evaluate all final models on the full SuperGLEBer benchmark. Table˜2 reports average performances across a subset of our models. In particular, we compare our models to established encoders: GBERT (Chan et al., 2020), GeBERTa (Dada et al., 2023), XLM-RoBERTa (Conneau et al., 2020; Goyal et al., 2021), mmBERT Marone et al. (2025) and EuroBERT (Boizard et al., 2025).
Our ModernGBERT consistently outperforms comparable and larger models. The 134M variant achieves an average score of 0.749, surpassing all similar-sized baselines, as well as LLäMmlein 1B (0.733), EuroBERT-2b (0.731) and LLäMmlein 7B (0.747). The ModernGBERT 1B variant achieves a new state-of-the-art average score across the entire SuperGLEBer of 0.808, outperforming GBERTLarge (0.768) by 4% and beating the seven times larger LLäMmlein2Vec 7B (0.787). It leads in three of four evaluation categories, including classification (0.812), NER (0.845), and QA (0.876). Only on sentence similarity (0.699), LLäMmlein2Vec 7B achieves better results. ModernGBERT scales well, with performance improving for larger model sizes, again suggesting that scaling ModernBERT-style encoders can leverage large monolingual corpora effectively. In the SuperGLEBer setting, adding context extension improved ModernGBERT’s average by 1.9% for the 134M model (from 0.730 to 0.749) and by 0.8% for the 1B variant (from 0.800 to 0.808). However, large improvements were not expected, as SuperGLEBer tasks do not make use of long contexts.
Adaptation via LLM2Vec yields consistent gains across models. Our first LLM2Vec tuning (on ext1, Section˜2.2) showed the most prominent positive effect, while the second fine-tune using the ext2 datasets showed only marginal increase, or even a decrease in performance. The same holds for a mixture of the two LLM2Vec adapters (ext1+2). Therefore, we report only ext1 results in this and all subsequent tables for simplicity. The LLäMmlein2Vec 7B achieves the strongest results among the LLM2Vec models (0.787). Conversion of LLäMmlein 120M, 1B, 7B improved the average score by 0.8%, 2.9%, and 4.0% respectively. This effect is especially pronounced in PAWSX, with scores increasing by up to 14.6% for LLäMmlein 7B and 6.7% for LLäMmlein 1B.
Comparing the LLäMmlein2Vec with the ModernGBERT family, we find that on similarly sized models, ModernGBERT always outperforms the transformed decoders by a large margin. Only the much larger LLäMmlein2Vec 7B approaches the performance of ModernGBERT 1B. This systematic comparison using identical datasets provides the first comprehensive analysis of MLM vs. LLM2Vec for encoder development.
| Model | Size | Avg. | Clstr | ReRnk | Retr |
| GBERTBase | 111M | 0.360 | 0.274 | 0.118 | 0.226 |
| GBERTBase f | 111M | 0.500 | 0.318 | 0.374 | 0.461 |
| GBERTLarge | 337M | 0.412 | 0.336 | 0.206 | 0.297 |
| GBERTLarge f | 337M | 0.521 | 0.334 | 0.389 | 0.493 |
| GeBERTaBase | 139M | 0.382 | 0.312 | 0.174 | 0.213 |
| GeBERTaBase f | 139M | 0.493 | 0.318 | 0.374 | 0.430 |
| GeBERTaLarge | 406M | 0.397 | 0.287 | 0.223 | 0.274 |
| GeBERTaLarge f | 406M | 0.494 | 0.311 | 0.374 | 0.432 |
| GeBERTaXLarge | 887M | 0.325 | 0.278 | 0.108 | 0.058 |
| GeBERTaXLarge f | 887M | 0.521 | 0.323 | 0.414 | 0.462 |
| XLM-RoBERTaBase | 279M | 0.248 | 0.173 | 0.024 | 0.008 |
| XLM-RoBERTaBase f | 279M | 0.403 | 0.247 | 0.247 | 0.299 |
| XLM-RoBERTaLarge | 561M | 0.264 | 0.172 | 0.048 | 0.026 |
| XLM-RoBERTaLarge f | 561M | 0.460 | 0.259 | 0.343 | 0.416 |
| XLM-RoBERTaXLarge | 3.48B | 0.301 | 0.225 | 0.090 | 0.142 |
| XLM-RoBERTaXLarge f | 3.48B | 0.479 | 0.342 | 0.362 | 0.407 |
| mmBERT-small | 142M | 0.289 | 0.163 | 0.091 | 0.075 |
| mmBERT-small f | 142M | 0.430 | 0.167 | 0.390 | 0.359 |
| mmBERT-base | 309M | 0.318 | 0.237 | 0.092 | 0.071 |
| mmBERT-base f | 309M | 0.475 | 0.214 | 0.428 | 0.406 |
| EuroBERT-210m | 212M | 0.293 | 0.240 | 0.070 | 0.118 |
| EuroBERT-210m f | 212M | 0.431 | 0.189 | 0.404 | 0.371 |
| EuroBERT-610m | 609M | 0.299 | 0.232 | 0.093 | 0.134 |
| EuroBERT-610m f | 609M | 0.419 | 0.288 | 0.263 | 0.352 |
| EuroBERT-2b | 2.11B | 0.230 | 0.186 | 0.060 | 0.065 |
| EuroBERT-2b f | 2.11B | 0.452 | 0.191 | 0.460 | 0.420 |
| LLäMmlein2Vec (ext1) | 120M | 0.315 | 0.261 | 0.139 | 0.224 |
| LLäMmlein2Vec (ext1)f | 120M | 0.471 | 0.308 | 0.325 | 0.425 |
| LLäMmlein2Vec (ext1) | 1B | 0.399 | 0.308 | 0.183 | 0.276 |
| LLäMmlein2Vec (ext1)f | 1B | 0.540 | 0.343 | 0.433 | 0.511 |
| LLäMmlein2Vec (ext1) | 7B | 0.376 | 0.249 | 0.169 | 0.266 |
| LLäMmlein2Vec (ext1)f | 7B | 0.557 | 0.339 | 0.477 | 0.522 |
| ModernGBERT | 134M | 0.383 | 0.293 | 0.139 | 0.241 |
| ModernGBERT f | 134M | 0.485 | 0.303 | 0.364 | 0.432 |
| ModernGBERT (ext1+2) | 134M | 0.376 | 0.296 | 0.120 | 0.213 |
| ModernGBERT (ext1+2) f | 134M | 0.501 | 0.312 | 0.404 | 0.446 |
| ModernGBERT | 1B | 0.374 | 0.318 | 0.097 | 0.199 |
| ModernGBERT f | 1B | 0.549 | 0.339 | 0.463 | 0.511 |
| ModernGBERT (ext1+2) | 1B | 0.366 | 0.307 | 0.088 | 0.191 |
| ModernGBERT (ext1+2) f | 1B | 0.551 | 0.338 | 0.459 | 0.512 |
Text Embedding
We evaluate models on the MTEB benchmark, which covers six task categories: classification, pair classification, clustering, reranking, retrieval, and short text similarity (STS) tasks. A summary of results is visualized in Table˜3. In general, supervised fine-tuning on mMARCO yields consistent improvements across all model types. While classification performance sometimes declines, substantial gains can be observed in other areas: 25% on average for reranking, 26% for retrieval and 25% for STS.
The best overall average performance is achieved by the fine-tuned LLäMmlein2Vec 7B (0.557), closely followed by the fine-tuned ModernGBERT 1B (0.551), despite the latter being significantly smaller. LLäMmlein2Vec models generally show strong performance after fine-tuning, particularly when trained with the extension dataset of the first phase (ext1). Using the second extension phase (ext2) or combining both adapters into the base model (ext1+2) again harms the performance. Interestingly, the latter shows the largest fine-tuning gains among the three variants. The ModernGBERT models perform competitively to similarly sized models. Before fine-tuning, ModernGBERT 1B (avg. 0.366) already outperforms most encoder-only models, such as GeBERTaXLarge (0.325), XLM-RoBERTaXLarge (0.301), and EuroBERT-2b (0.230) but not GBERTLarge (0.412). However, after fine-tuning, it demonstrates clear superiority among native encoder-only models by at least 3% on the average score. As with our observations on the SuperGLEBer benchmark, ModernGBERT’s context extension did not show significant improvements here. Comparing ModernGBERT and LLäMmlein2Vec, we find that before fine-tuning, the LLäMmlein2Vec 1B and 7B models produce better representations than ModernGBERT 1B. However, after fine-tuning, ModernGBERT 1B surpasses the 1B variant of LLäMmlein2Vec on average and closely aligns with the larger 7B model.
Long-Context Understanding
Table˜4 reports results on our new German QA-NIAH benchmark . We evaluate subsets of short (), medium ( to ), and long ( to ) sequences, focusing on LLMs supporting up to tokens: ModernGBERT, the encoder-converted LLäMmlein2Vec, as well as their original decoder counterparts. Notably, LLäMmlein models were pre-trained with a maximum context of tokens and only LLäMmlein2Vec was post-trained with a maximum context length of . ModernGBERT 1B demonstrates strong long-context performance across all lengths, outperforming all encoders. The first extension phase during training approximately tripled accuracy, while the final High Quality extension slightly reduced performance, especially for the 134M variant. Regarding LLM2Vec, a sufficiently long conversion improved long-context understanding. Conversion of LLäMmlein 120M and 1B decoders (with native context length of ) improved accuracy by factor 1.3 resp. 2, both not as pronounced in comparison to the ModernGBERT encoders. For LLäMmlein2Vec 7B however (with LLM2Vec training on approximately half of our ext1 dataset), it decreased by 51%, with no correct answers on haystacks of tokens. Given the intensive compute requirements, we did not explore further optimizations regarding context extension of the LLäMmlein2Vec 7B model.
| Model | Size | tok. | to tok. | to tok. | Avg. |
| LLäMmlein | 120M | 0.286 | 0.124 | 0.049 | 0.091 |
| LLäMmlein | 1B | 0.517 | 0.230 | 0.088 | 0.165 |
| LLäMmlein | 7B | 0.529 | 0.310 | 0.122 | 0.216 |
| LLäMmlein2Vec (ext1) | 120M | 0.315 | 0.206 | 0.044 | 0.120 |
| LLäMmlein2Vec (ext1) | 1B | 0.588 | 0.448 | 0.232 | 0.333 |
| LLäMmlein2Vec (ext1) | 7B | 0.597 | 0.207 | 0.000 | 0.111 |
| ModernGBERT | 134M | 0.552 | 0.168 | 0.013 | 0.105 |
| ModernGBERT (ext1) | 134M | 0.536 | 0.410 | 0.238 | 0.323 |
| ModernGBERT (ext1+2) | 134M | 0.540 | 0.393 | 0.201 | 9.296 |
| ModernGBERT | 1B | 0.556 | 0.233 | 0.023 | 0.136 |
| ModernGBERT (ext1) | 1B | 0.617 | 0.506 | 0.406 | 0.457 |
| ModernGBERT (ext1+2) | 1B | 0.526 | 0.383 | 0.451 | 0.451 |
5.3. Inference Efficiency
| Long | |||
| Model | Size | Fixed Length | Variable Length |
| LLäMmlein2Vec | 120M | p m 0.14 | p m 0.35 |
| LLäMmlein2Vec | 1B | p m 0.12 | p m 0.30 |
| LLäMmlein2Vec | 7B | p m 0.19 | p m 0.41 |
| ModernGBERT | 134M | ||
| ModernGBERT | 1B | ||
We evaluate inference efficiency across varying sequence lengths using four synthetic datasets, each containing documents composed of random tokens. Following Warner et al. (2024), two datasets use fixed-length sequences ( and tokens), while the other two sample sequence lengths from normal distributions (either mean 256, variance 64; or mean , variance ) to better simulate real-world conditions. Our ModernGBERT models adopt ModernBERT’s unpadding approach: padding tokens are removed and sequences in a batch are concatenated, allowing Flash Attention to handle variable-length attention masks. The computational equivalence is facilitated by carefully crafting an appropriate attention mask. All other models rely on conventional padding. Among smaller models (134M resp. 120M), ModernGBERT and LLäMmlein2Vec achieve comparable efficiency on fixed-length data, both only surpassed by GBERTBase and XLM-RoBERTaBase in terms of efficiency on short sequences. For 1B variants, ModernGBERT consistently outperforms LLäMmlein2Vec 1B and 7B variations in inference speed, likely due to its architectural decisions optimized for efficiency, such as ensuring that weight matrices have dimension of multiples of 64, and are divisible into 128 256 block for efficient tiling on the GPU. Gains are most pronounced for variable-length datasets, where ModernGBERT’s unpadding yields clear benefits (see Table˜5): the 134M ModernGBERT is the most efficient model on variable length, and the 1B variant substantially outpaces its LLäMmlein2Vec counterpart. Notably, ModernGBERT 1B matches the MTEB performance of LLäMmlein2Vec 7B (see Table˜3) while being ten times faster on long-context documents.
6. Related Work
Next-Generation Encoders
Several recent efforts have extended ModernBERT to other languages and domains, such as French (Antoun et al., 2024) and Japanese (Sugiura et al., 2025). Recently, mmBERT was created - a ModernBERT model trained on 3T tokens of multilingual text covering over languages (Marone et al., 2025).
Concurrent to our work, several alternative encoder architectures have been proposed. Breton et al. (2025) introduced NeoBERT, a 250M parameter English encoder incorporating similar architectural innovations like ModernBERT, but scaling up layers rather than hidden dimension, switching from GeLU to SwiGLU activation, and using a modified training scheme (Cosine scheduler, reduced masking). Their model surpasses ModernBERT-large on GLUE and MTEB with 100M fewer parameters, although its scalability with model size remains unexplored. Likewise, Boizard et al. (2025) recently presented EuroBERT (210M, 610M, 2.1B), a multilingual encoder family featuring architectural changes similar to those of ModernBERT, but retaining some architectural details (RMSNorm layer normalization, SiLU activation function, Llama-style tokenizer) from the Llama family, resembling our LLäMmlein2Vec architecture.
Antoun et al. (2025) compared French ModernBERT and DeBERTaV3, finding DeBERTaV3 to be superior on downstream tasks but significantly slower in training and inference.
Tuning decoder-only LLMs into Encoders
Due to the lack of new encoder models (Weller et al., 2025), works like LLM2Vec (Section˜3.2) have proposed strategies for turning decoder-only models into encoders (Section˜3.2). Few works have investigated converting decoder-only LLMs into encoders, besides LLM2Vec (Section˜3.2). Recent studies predominantly address either distilling text embedders (Li and Li, 2024; Lee et al., 2025, 2024; Ma et al., 2025) or fine-tuning LLMs as bidirectional encoders for specific tasks (Li et al., 2023; Dukić and Snajder, 2024), with evaluation primarily focused on English or multilingual settings.
Khosla et al. (2025) introduced MAGNET, a method for converting decoder-only models into foundational encoders, similarly to LLM2Vec. Unlike LLM2Vec, which enables bidirectional attention and masked next-token prediction, MAGNET uses a hybrid of bidirectional and causal attention and adds a missing-span generation objective, aiming for a more general pre-training signal. Concurrent to our work, the Ettin Suite (Weller et al., 2025) provides a systematic comparison of encoder-only and decoder-only architectures trained on identical data and recipes. Notably, they also experiment with objective switching, continuing training on 50B tokens, and directly comparing downstream task performance.
7. Conclusion
In this work we introduce two encoder-only families from scratch ModernGBERT and LLäMmlein2Vec, contributing: a) the first systematic MLM vs. LLM2Vec comparison for encoders using identical datasets, b) a novel QA-NIAH evaluation adapted for encoder models, and c) the first combination of LLM2Vec with context extension training.
Our main findings show that the proposed ModernGBERT family, especially the 1B variant, sets a new state-of-the-art for German encoders, outperforming previous models while remaining suitable for practical deployment as a drop-in replacement for GBERT, capable of handling sequences of up to tokens. Our learning dynamics analysis confirms that larger encoder architectures can effectively exploit terabyte-scale German monolingual corpora, with performance consistently improving with increased model size and data. A comparison of ModernGBERT and LLäMmlein2Vec (derived from LLäMmlein) shows that dedicated encoder training yields superior results, justifying its computational expense when parameter efficiency is essential. At the same time, LLM2Vec provides a resource-efficient alternative for scenarios where training a full encoder from scratch is too expensive. Notably, in some use cases, we find that fine-tuning a decoder with a bidirectional attention mask already delivers substantial gains. These trends suggest that even larger encoder models could yield further gains, which we leave to future work. By releasing ModernGBERT and LLäMmlein2Vec, along with full training transparency, intermediate checkpoints, detailed documentation, and accompanying resources, we aim to facilitate further development and understanding within the German LLM community.
Limitations
Despite the ModernGBERT and LLäMmlein2Vec models being a notable advancement in the German NLP landscape, several limitations persist: 1) Monolingual focus.Although the focus on German is a strength for this specific context, ModernGBERT is unable to utilize multilingual contexts or perform cross-lingual tasks, hindering applicability in some scenarios. 2) Limited coding capabilities.High-quality German resources for coding are rare, and no code is included in the training dataset. This restricts its capabilities in code retrieval applications. 3) Evaluation scope.While we rigorously evaluated our models on the German SuperGLEBer and MTEB benchmarks, these benchmarks are limited in their domain, and other domains such as literature, medical domains, or technical subjects were not tested. Furthermore, our benchmarks do not strictly probe for “German factual knowledge”, for instance, knowledge about German geography, or common German TV shows. 4) No custom tokenizer.We utilized the original BERT-style GBERT tokenizer due to its availability and persistent usage. However, we did not invest in developing a custom tokenizer, like the BPE-style OLMo tokenizer used in ModernBERT. Consequently, ModernGBERT’s tokenizer cannot, e.g., differentiate between various whitespace characters or encode emoji. 5) Evaluation of long-context understanding.Due to the absence of high-quality native German evaluation datasets, we had to rely on non-natural QA-NIAH sequences, only broadly testing for long-context understanding. Contrast this with English benchmarks such as Bench-MC (Zhang et al., 2024) or LongBench-v2 (Bai et al., 2025), which include full novels along with questions that require attention to much information scattered throughout the novel. In future work, we plan on developing a dedicated high-quality non-synthetic German long-context evaluation benchmark.
8. Bibliographical References
- Antoun et al. (2024) Wissam Antoun, Francis Kulumba, Rian Touchent, Éric de la Clergerie, Benoît Sagot, and Djamé Seddah. 2024. Camembert 2.0: A smarter french language model aged to perfection.
- Antoun et al. (2025) Wissam Antoun, Benoît Sagot, and Djamé Seddah. 2025. ModernBERT or DeBERTaV3? examining architecture and data influence on transformer encoder models performance. ArXiv:2504.08716.
- Bai et al. (2025) Yushi Bai, Shangqing Tu, Jiajie Zhang, Hao Peng, Xiaozhi Wang, Xin Lv, Shulin Cao, Jiazheng Xu, Lei Hou, Yuxiao Dong, Jie Tang, and Juanzi Li. 2025. LongBench v2: Towards deeper understanding and reasoning on realistic long-context multitasks. ArXiv:2412.15204.
- BehnamGhader et al. (2024) Parishad BehnamGhader, Vaibhav Adlakha, Marius Mosbach, Dzmitry Bahdanau, Nicolas Chapados, and Siva Reddy. 2024. LLM2Vec: Large language models are secretly powerful text encoders. In First Conference on Language Modeling, Philadelphia, PA, USA.
- Biderman et al. (2023) Stella Biderman, Hailey Schoelkopf, Quentin Gregory Anthony, Herbie Bradley, Kyle O’Brien, Eric Hallahan, Mohammad Aflah Khan, Shivanshu Purohit, Usvsn Sai Prashanth, Edward Raff, Aviya Skowron, Lintang Sutawika, and Oskar Van Der Wal. 2023. Pythia: A suite for analyzing large language models across training and scaling. In Proceedings of the 40th International Conference on Machine Learning, volume 202 of Proceedings of Machine Learning Research, pages 2397–2430. PMLR.
- Boizard et al. (2025) Nicolas Boizard, Hippolyte Gisserot-Boukhlef, Duarte Miguel Alves, Andre Martins, Ayoub Hammal, Caio Corro, CELINE HUDELOT, Emmanuel Malherbe, Etienne Malaboeuf, Fanny Jourdan, Gabriel Hautreux, João Alves, Kevin El Haddad, Manuel Faysse, Maxime Peyrard, Nuno M Guerreiro, Patrick Fernandes, Ricardo Rei, and Pierre Colombo. 2025. EuroBERT: Scaling multilingual encoders for european languages. In Second Conference on Language Modeling.
- Bonifacio et al. (2022) Luiz Bonifacio, Vitor Jeronymo, Hugo Queiroz Abonizio, Israel Campiotti, Marzieh Fadaee, Roberto Lotufo, and Rodrigo Nogueira. 2022. mMARCO: A multilingual version of the MS MARCO passage ranking dataset. ArXiv:2108.13897.
- Bourdois (2025) Loïck Bourdois. 2025. Model statistics of the 50 most downloaded entities on hugging face.
- Breton et al. (2025) Lola Le Breton, Quentin Fournier, Mariam El Mezouar, and Sarath Chandar. 2025. NeoBERT: A next-generation BERT. ArXiv:2502.19587.
- Chan et al. (2020) Branden Chan, Stefan Schweter, and Timo Möller. 2020. German‘s next language model. In Proceedings of the 28th International Conference on Computational Linguistics, pages 6788–6796, Barcelona, Spain (Online). International Committee on Computational Linguistics.
- Conneau et al. (2020) Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume Wenzek, Francisco Guzmán, Edouard Grave, Myle Ott, Luke Zettlemoyer, and Veselin Stoyanov. 2020. Unsupervised cross-lingual representation learning at scale. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 8440–8451, Online. Association for Computational Linguistics.
- Conneau et al. (2018) Alexis Conneau, Ruty Rinott, Guillaume Lample, Adina Williams, Samuel Bowman, Holger Schwenk, and Veselin Stoyanov. 2018. XNLI: Evaluating cross-lingual sentence representations. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing, pages 2475–2485, Brussels, Belgium. Association for Computational Linguistics.
- Dada et al. (2023) Amin Dada, Aokun Chen, Cheng Peng, Kaleb Smith, Ahmad Idrissi-Yaghir, Constantin Seibold, Jianning Li, Lars Heiliger, Christoph Friedrich, Daniel Truhn, Jan Egger, Jiang Bian, Jens Kleesiek, and Yonghui Wu. 2023. On the impact of cross-domain data on German language models. In Findings of the Association for Computational Linguistics: EMNLP 2023, pages 13801–13813, Singapore. Association for Computational Linguistics.
- Dettmers et al. (2023) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer. 2023. QLoRA: Efficient finetuning of quantized llms. ArXiv:2305.14314.
- Dukić and Snajder (2024) David Dukić and Jan Snajder. 2024. Looking right is sometimes right: Investigating the capabilities of decoder-only LLMs for sequence labeling. In Findings of the Association for Computational Linguistics: ACL 2024, pages 14168–14181, Bangkok, Thailand. Association for Computational Linguistics.
- Enevoldsen et al. (2025) Kenneth Enevoldsen, Isaac Chung, Imene Kerboua, Márton Kardos, Ashwin Mathur, David Stap, Jay Gala, Wissam Siblini, Dominik Krzemiński, Genta Indra Winata, Saba Sturua, Saiteja Utpala, Mathieu Ciancone, Marion Schaeffer, Gabriel Sequeira, Diganta Misra, Shreeya Dhakal, Jonathan Rystrøm, Roman Solomatin, Ömer Çağatan, Akash Kundu, Martin Bernstorff, Shitao Xiao, Akshita Sukhlecha, Bhavish Pahwa, Rafał Poświata, Kranthi Kiran GV, Shawon Ashraf, Daniel Auras, Björn Plüster, Jan Philipp Harries, Loïc Magne, Isabelle Mohr, Mariya Hendriksen, Dawei Zhu, Hippolyte Gisserot-Boukhlef, Tom Aarsen, Jan Kostkan, Konrad Wojtasik, Taemin Lee, Marek Šuppa, Crystina Zhang, Roberta Rocca, Mohammed Hamdy, Andrianos Michail, John Yang, Manuel Faysse, Aleksei Vatolin, Nandan Thakur, Manan Dey, Dipam Vasani, Pranjal Chitale, Simone Tedeschi, Nguyen Tai, Artem Snegirev, Michael Günther, Mengzhou Xia, Weijia Shi, Xing Han Lù, Jordan Clive, Gayatri Krishnakumar, Anna Maksimova, Silvan Wehrli, Maria Tikhonova, Henil Panchal, Aleksandr Abramov, Malte Ostendorff, Zheng Liu, Simon Clematide, Lester James Miranda, Alena Fenogenova, Guangyu Song, Ruqiya Bin Safi, Wen-Ding Li, Alessia Borghini, Federico Cassano, Hongjin Su, Jimmy Lin, Howard Yen, Lasse Hansen, Sara Hooker, Chenghao Xiao, Vaibhav Adlakha, Orion Weller, Siva Reddy, and Niklas Muennighoff. 2025. MMTEB: Massive multilingual text embedding benchmark. ArXiv:2502.13595.
- Faruqui and Padó (2010) Manaal Faruqui and Sebastian Padó. 2010. Training and evaluating a german named entity recognizer with semantic generalization. In Conference on Natural Language Processing.
- Gao et al. (2025) Tianyu Gao, Alexander Wettig, Howard Yen, and Danqi Chen. 2025. How to train long-context language models (effectively). ArXiv:2410.02660.
- Goyal et al. (2021) Naman Goyal, Jingfei Du, Myle Ott, Giri Anantharaman, and Alexis Conneau. 2021. Larger-scale transformers for multilingual masked language modeling. In Proceedings of the 6th Workshop on Representation Learning for NLP (RepL4NLP-2021), pages 29–33, Online. Association for Computational Linguistics.
- Henrich et al. (2012) Verena Henrich, Erhard Hinrichs, and Tatiana Vodolazova. 2012. WebCAGe – a web-harvested corpus annotated with GermaNet senses. In Proceedings of the 13th Conference of the European Chapter of the Association for Computational Linguistics, pages 387–396, Avignon, France. Association for Computational Linguistics.
- Hsieh et al. (2024) Cheng-Ping Hsieh, Simeng Sun, Samuel Kriman, Shantanu Acharya, Dima Rekesh, Fei Jia, Yang Zhang, and Boris Ginsburg. 2024. RULER: What’s the real context size of your long-context language models? ArXiv:2404.06654.
- Ivgi et al. (2023) Maor Ivgi, Uri Shaham, and Jonathan Berant. 2023. Efficient long-text understanding with short-text models. Transactions of the Association for Computational Linguistics, 11:284–299.
- Khosla et al. (2025) Savya Khosla, Aditi Tiwari, Kushal Kafle, Simon Jenni, Handong Zhao, John Collomosse, and Jing Shi. 2025. MAGNET: Augmenting generative decoders with representation learning and infilling capabilities. ArXiv:2501.08648.
- Lee et al. (2025) Chankyu Lee, Rajarshi Roy, Mengyao Xu, Jonathan Raiman, Mohammad Shoeybi, Bryan Catanzaro, and Wei Ping. 2025. NV-embed: Improved techniques for training llms as generalist embedding models. ArXiv:2405.17428.
- Lee et al. (2024) Jinhyuk Lee, Zhuyun Dai, Xiaoqi Ren, Blair Chen, Daniel Cer, Jeremy R. Cole, Kai Hui, Michael Boratko, Rajvi Kapadia, Wen Ding, Yi Luan, Sai Meher Karthik Duddu, Gustavo Hernandez Abrego, Weiqiang Shi, Nithi Gupta, Aditya Kusupati, Prateek Jain, Siddhartha Reddy Jonnalagadda, Ming-Wei Chang, and Iftekhar Naim. 2024. Gecko: Versatile text embeddings distilled from large language models. ArXiv:2403.20327.
- Li and Li (2024) Xianming Li and Jing Li. 2024. BeLLM: Backward dependency enhanced large language model for sentence embeddings. In Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 792–804, Mexico City, Mexico. Association for Computational Linguistics.
- Li et al. (2023) Zongxi Li, Xianming Li, Yuzhang Liu, Haoran Xie, Jing Li, Fu lee Wang, Qing Li, and Xiaoqin Zhong. 2023. Label supervised LLaMA finetuning. ArXiv:2310.01208.
- Liang et al. (2020) Yaobo Liang, Nan Duan, Yeyun Gong, Ning Wu, Fenfei Guo, Weizhen Qi, Ming Gong, Linjun Shou, Daxin Jiang, Guihong Cao, Xiaodong Fan, Ruofei Zhang, Rahul Agrawal, Edward Cui, Sining Wei, Taroon Bharti, Ying Qiao, Jiun-Hung Chen, Winnie Wu, Shuguang Liu, Fan Yang, Daniel Campos, Rangan Majumder, and Ming Zhou. 2020. XGLUE: A new benchmark dataset for cross-lingual pre-training, understanding and generation. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 6008–6018, Online. Association for Computational Linguistics.
- Ma et al. (2025) Xueguang Ma, Xi Victoria Lin, Barlas Oguz, Jimmy Lin, Wen tau Yih, and Xilun Chen. 2025. DRAMA: Diverse augmentation from large language models to smaller dense retrievers. ArXiv:2502.18460.
- Marone et al. (2025) Marc Marone, Orion Weller, William Fleshman, Eugene Yang, Dawn Lawrie, and Benjamin Van Durme. 2025. mmbert: A modern multilingual encoder with annealed language learning.
- Möller et al. (2021) Timo Möller, Julian Risch, and Malte Pietsch. 2021. GermanQuAD and GermanDPR: Improving non-English question answering and passage retrieval. In Proceedings of the 3rd Workshop on Machine Reading for Question Answering, pages 42–50, Punta Cana, Dominican Republic. Association for Computational Linguistics.
- Muennighoff et al. (2023) Niklas Muennighoff, Nouamane Tazi, Loïc Magne, and Nils Reimers. 2023. MTEB: Massive text embedding benchmark. ArXiv:2210.07316.
- Penedo et al. (2024) Guilherme Penedo, Hynek Kydlíček, Vinko Sabolčec, Bettina Messmer, Negar Foroutan, Martin Jaggi, Leandro von Werra, and Thomas Wolf. 2024. Fineweb2: A sparkling update with 1000s of languages. Huggingface:HuggingFaceFW/fineweb-2.
- Pfister and Hotho (2024) Jan Pfister and Andreas Hotho. 2024. SuperGLEBer: German language understanding evaluation benchmark. In Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 7904–7923, Mexico City, Mexico. Association for Computational Linguistics.
- Pfister et al. (2024) Jan Pfister, Julia Wunderle, and Andreas Hotho. 2024. LLäMmlein: Compact and competitive german-only language models from scratch. ArXiv:2411.11171.
- Reimers and Gurevych (2019) Nils Reimers and Iryna Gurevych. 2019. Sentence-BERT: Sentence embeddings using Siamese BERT-networks. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pages 3982–3992, Hong Kong, China. Association for Computational Linguistics.
- Risch et al. (2021) Julian Risch, Anke Stoll, Lena Wilms, and Michael Wiegand. 2021. Overview of the GermEval 2021 shared task on the identification of toxic, engaging, and fact-claiming comments. In Proceedings of the GermEval 2021 Shared Task on the Identification of Toxic, Engaging, and Fact-Claiming Comments, pages 1–12, Duesseldorf, Germany. Association for Computational Linguistics.
- Sugiura et al. (2025) Issa Sugiura, Kouta Nakayama, and Yusuke Oda. 2025. llm-jp-modernbert: A ModernBERT model trained on a large-scale japanese corpus with long context length. ArXiv:2504.15544.
- Warner et al. (2024) Benjamin Warner, Antoine Chaffin, Benjamin Clavié, Orion Weller, Oskar Hallström, Said Taghadouini, Alexis Gallagher, Raja Biswas, Faisal Ladhak, Tom Aarsen, Nathan Cooper, Griffin Adams, Jeremy Howard, and Iacopo Poli. 2024. Smarter, better, faster, longer: A modern bidirectional encoder for fast, memory efficient, and long context finetuning and inference. ArXiv:2412.13663.
- Weber et al. (2024) Maurice Weber, Daniel Y. Fu, Quentin Anthony, Yonatan Oren, Shane Adams, Anton Alexandrov, Xiaozhong Lyu, Huu Nguyen, Xiaozhe Yao, Virginia Adams, Ben Athiwaratkun, Rahul Chalamala, Kezhen Chen, Max Ryabinin, Tri Dao, Percy Liang, Christopher Ré, Irina Rish, and Ce Zhang. 2024. RedPajama: an open dataset for training large language models. In Advances in Neural Information Processing Systems, volume 37, pages 116462–116492. Curran Associates, Inc.
- Wehrli et al. (2023) Silvan Wehrli, Bert Arnrich, and Christopher Irrgang. 2023. German text embedding clustering benchmark. In Proceedings of the 19th Conference on Natural Language Processing (KONVENS 2023), pages 187–201, Ingolstadt, Germany. Association for Computational Lingustics.
- Weller et al. (2025) Orion Weller, Kathryn Ricci, Marc Marone, Antoine Chaffin, Dawn Lawrie, and Benjamin Van Durme. 2025. Seq vs seq: An open suite of paired encoders and decoders.
- Wojatzki et al. (2017) Michael Wojatzki, Eugen Ruppert, Sarah Holschneider, Torsten Zesch, and Chris Biemann. 2017. Germeval 2017: Shared task on aspect-based sentiment in social media customer feedback. In Proceedings of the GermEval 2017 – Shared Task on Aspect-based Sentiment in Social Media Customer Feedback, pages 1–12, Berlin, Germany.
- Zhang et al. (2024) Xinrong Zhang, Yingfa Chen, Shengding Hu, Zihang Xu, Junhao Chen, Moo Hao, Xu Han, Zhen Thai, Shuo Wang, Zhiyuan Liu, and Maosong Sun. 2024. Bench: Extending long context evaluation beyond 100K tokens. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 15262–15277, Bangkok, Thailand. Association for Computational Linguistics.